OpenStudy (anonymous):

3.Angelo took a group of people to the movies and spent $80 on tickets. Adult tickets cost $10 and child tickets cost $6. The equation to represent his cost is 10A + 6C = $80. If there were 5 children in the group, how many adult tickets did he purchase? Show your work and use a complete sentence to explain your solution in the context of the original problem.

OpenStudy (anonymous):

10A + 6C = $80 1) Substitute 5 for C. 10A+6(5)=80 2)Multiply 6*5. 10A+30=80 3)Subtract 30 from both sides 10A=50 4)Divide both sides by 10. A=5 Angelo purchased 5 adult tickets.

OpenStudy (anonymous):

thank you very much hun :)

OpenStudy (anonymous):

You're welcome ^u^